you have found the following historical information for the Daniela company
year 1 year 2 year 3 year 4
stock
price. 46.02 60.43 66.09 6301
EPS. 2.45 2.51 2.68 2.67
Earnings are expected to grow at 12 percent for the next year stock
using the compqnys historical average PE as a benchmark what is the tatget stoxk price in one year?
Answer #1

Step 1 - Calculation of P/E

Year 1 Year 2 Year 3 Year 4
Stock price 46.02 60.43 66.09 63.01
EPS 2.45 2.51 2.68 2.67
P/E 18.78 24.08 24.66 23.6

Step 2 - Calculation of average P/E

(18.78+24.08+24.66+23.66) / 4=91.12 / 4 = 22.78

Step 3 - Calculation of year 5 EPS

2.67*1.12 = 2.99

Step 4 - Calculation of Year 5 stock price

2.99 * 22.78 = 68.12